Acetylene, C2H2 , is used in welding torches. It releases a lot of energy when burned in oxygen. The combustion of one gram of acetylene releases 48.2 kJ. A 0.750-gram sample of acetylene is burned in a bomb calorimeter (heat capacity = 1.117kJ/OC) that contains 800.0 g of water. The final temperature of the bomb and water after combustion is 35.2OC. What is the initial temperature of the bomb and water?


1
Expert's answer
2021-10-28T02:58:57-0400

Heat released by combustion of 1 gram of acetylene = 48.2 kJ

So for combustion of 0.750 grams,

Q = 0.75 × 48.2

Q = 10.75 × 48.2​

Q = 36.15 kJ

Q = C × ΔT

C = Heat capacity = 1.117 kJ/°C

ΔT = Temperature difference = T2 - T1

T2 = Final temperature = 35.2°C

T1 = Initial temperature

Substitute the values,

ΔT = 32.36 °C

32.36℃ = T2 - T1

T1 = 35.2 - 32.36

T1 = 2.84 °C